# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 7.8 Dual exceptional collections
We know this already for $M,N\in{\mathrm{rep}}_{\mathrm{fr}}(\mathbb{B})$, cf. Remark 6.8. By Proposition 8.1 the results follow. ∎
8.2 The category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ as a $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$-linear category
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. The restriction functor ${\mathrm{res}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(B))$ is $t$-exact. Its right adjoint is the induction functor ${\mathrm{Rind}}_{B}^{G}$ and the fact that ${\mathrm{res}}_{B}^{G}$ gives a full faithful embedding is a consequence of the Kempf vanishing theorem, [10, Theorem 1.2], [29, B.3] and Section 14.2: it implies that ${\mathrm{Rind}}_{B}^{G}{\mathrm{res}}_{B}^{G}={\sf id}_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))}$.
@@
The functor ${\mathrm{Rind}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))$ restricts to a functor ${\mathrm{Rind}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$.
@@
Set $X=G/B$, $Y={\sf pt}$, $f:X\rightarrow Y$, $H=G$ in [37, Proposition A.12]. ∎
@@
Proof.
Put $X=G/B,H=G$ in [37, Appendix, Corollary A.5]
[…]
Let ${\sf Inv}^{G}$ denote the derived functor of invariants $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\sf k}\textrm{-}\mathrm{Mod})$.
Proposition 8.6.
Let ${\sf F},{\sf G}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{rep}}(B))$ and $\mathcal{F}=\mathcal{L}({\sf F})$, ${\mathcal{G}}=\mathcal{L}({\sf G})\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}_{G}(G/B)$ be the associated equivariant complexes of coherent sheaves. There there is a canonical functorial isomorphism
[…]
Proof.
Put $X=G/B$ and $H=G$ in [37, Appendix, Proposition A.6]
[…]
8.3 $G$-linear semiorthogonal decompositions
Definition 19.
A triangulated category $\mathcal{D}$ is called $G$-linear if $\mathcal{D}$ is equipped with a monoidal action of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$, i.e. there is a bifunctor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\times\mathcal{D}\rightarrow\mathcal{D}$. Compare [31, Section 2.7].
Our main concern is the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$ which becomes a $G$-linear triangulated category under the restriction functor ${\mathrm{res}}_{B}^{G}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(B))$. Later we will also need $P$, so let us use $P$ instead of $B$, having in mind $P=B$ as an important case.
Definition 20.
A $G$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ is a collection $\langle{\sf A}_{1},\dots\sf A_{k}\rangle$ of full triangulated $G$-linear subcategories such that the collection $\langle{\sf A}_{1},\dots\sf A_{k}\rangle$ is a semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ in the sense of Definition 14.
The following proposition is just a variation of [33, Lemma 2.7]:
Proposition 8.7.
A pair of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$-linear subcategories ${\sf A},{\sf B}\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ is semiorthogonal (i.e.  ${\sf A}\subset{\sf B}^{\perp}$) if and only if the equality ${\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{k}M)=0$ holds for all $M\in{\sf A}$, $N\in{\sf B}$.
Proof.
($\Leftarrow$). Assume given $M\in{\sf A}$, $N\in{\sf B}$ and assume ${\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{k}M)=0$. Then
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}(N,M)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}({\sf k},N^{*}\otimes M)=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}({\sf k},{\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{k}M))=0.$
($\Rightarrow$). Let $M\in{\sf A}$, $N\in{\sf B}$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}(N,M)=0$ for all such $M$, $N$. Let be $L$ an arbitrary object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. Then
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(L,{\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{{\sf k}}M))=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}({\mathrm{res}}_{P}^{G}L,N^{\ast}\otimes_{{\sf k}}M)$
and
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}({\mathrm{res}}_{P}^{G}L,N^{\ast}\otimes_{{\sf k}}M)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}(N\otimes_{{\sf k}}{\mathrm{res}}_{P}^{G}L,M)=0;$
@@ 8.3 $G$-linear semiorthogonal decompositions
the latter equality holds since the subcategory $\sf B$ is $G$-linear and hence stable under tensoring with objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. It follows that $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}(L,{\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{{\sf k}}M))=0$ for an arbitrary object $L\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$; hence, ${\mathrm{Rind}}_{P}^{G}(N^{\ast}\otimes_{{\sf k}}M)=0$. ∎
@@
Let us simplify notation and write $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))}$ as $\mathop{\mathrm{RHom}}\nolimits_{P}$, and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))}$ as $\mathop{\mathrm{RHom}}\nolimits_{G}$.
Definition 21.
@@
A functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(P))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(P))$ is called $G$-linear if for all $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(P))$, $N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ there are given bifunctorial isomorphisms
$\Phi({\mathrm{res}}_{P}^{G}(N)\otimes M)={\mathrm{res}}_{P}^{G}(N)\otimes\Phi(M).$
@@
    Assume given a fully faithful $G$-linear functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. Then $\Phi$ is isomorphic to the functor $\Phi_{E}(-)=(-)\otimes E$ where $E\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ such that $\mathop{\mathrm{RHom}}\nolimits_{P}(E,E)={\sf k}$.
@@
    Each object $E\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$ such that ${\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E)={\sf k}$ gives a fully faithful $G$-linear functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$.
@@
    Under the assumptions of (1), the object $E$ satisfies ${\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E)={\sf k}$.
@@
(2) For $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$ we have
$\mathop{\mathrm{RHom}}\nolimits_{P}(M\otimes E,N\otimes E)=\mathop{\mathrm{RHom}}\nolimits_{P}(M\otimes N^{\ast},E^{\ast}\otimes E)=$
$\mathop{\mathrm{RHom}}\nolimits_{G}(M\otimes N^{\ast},{\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E))=\mathop{\mathrm{RHom}}\nolimits_{G}(M\otimes N^{\ast},{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{G}(M,N).$
@@
$\mathop{\mathrm{RHom}}\nolimits_{G}(M,{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{P}(M\otimes E,E)=$
$\mathop{\mathrm{RHom}}\nolimits_{P}(M,E^{\ast}\otimes E)=\mathop{\mathrm{RHom}}\nolimits_{G}(M,{\mathrm{Rind}}_{P}^{G}(E^{\ast}\otimes E)).$
for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(G))$. By the Yoneda Lemma it follows that ${\mathrm{Rind}}_{P}^{G}(X_{v}^{\ast}\otimes E)={\sf k}$. ∎
@@
See Section 6.1 for the precise relation between the two vanishing conditions ${\rm R}^{i}{\mathrm{ind}}_{B}^{G}(E^{\ast}\otimes E)=0$ for $i>0$ and $\mathop{\mathrm{RHom}}\nolimits^{i}_{B}(E,E)=0$ for $i>0$.
8.4 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(G/P))$
Definition 22.
@@
The set of line bundles $\langle\mathcal{L}_{\lambda}[n],\lambda\in X(T)_{+},n\in\mathbb{Z}\rangle$ is a generating set for $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(G/P))$.
@@
The smallest thick full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P)$ containing $\langle\mathcal{L}_{\lambda}[n],\lambda\in X(T)_{+},n\in\mathbb{Z}\rangle$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P)$.

### v3

@@ 7.8 Dual exceptional collections
We know this already for $M,N\in{\mathrm{rep}}_{\mathrm{fr}}(\mathbb{G})$, cf. Remark 6.8. By Proposition 8.1 the results follow. ∎
8.2 The category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ as a $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$-linear category
Let ${\sf k}$ be a field or ${\mathbb{Z}}$. The restriction functor ${\mathrm{res}}_{\mathbf{B}}^{\mathbf{G}}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}({\mathbf{B}}))$ is $t$-exact. Its right adjoint is the induction functor ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}$ and the fact that ${\mathrm{res}}_{\mathbf{B}}^{\mathbf{G}}$ gives a full faithful embedding is a consequence of the Generalized Tensor Identity [32, I Proposition 4.8] and the Kempf vanishing theorem, [13, Theorem 1.2], [32, B.3] or Section 14.2: it implies that ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}{\mathrm{res}}_{\mathbf{B}}^{\mathbf{G}}={\sf id}_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{G}))}$.
@@
The functor ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{G}))$ restricts to a functor ${\mathrm{Rind}}_{\mathbf{B}}^{\mathbf{G}}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$.
@@
Set $X=\mathbf{G}/\mathbf{B}$, $Y={\sf pt}$, $f:X\rightarrow Y$, $H=\mathbf{G}$ in [40, Proposition A.12]. ∎
It follows from the above that $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ identifies with a right admissible subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$. Both categories are monoidal and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}({\mathbf{B}}))$ is a $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$-linear triangulated category. Denote $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}_{\mathbf{G}}(\mathbf{G}/\mathbf{B}):=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathsf{QCoh}}^{\mathbf{G}}(\mathbf{G}/\mathbf{B}))$ and $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{\mathbf{G}}(\mathbf{G}/\mathbf{B}):=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\sf Coh}^{\mathbf{G}}(\mathbf{G}/\mathbf{B}))$.
Proposition 8.5.
Let ${\sf F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ and $\mathcal{F}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})$ be the associated complex of equivariant sheaves on $\mathbf{G}/\mathbf{B}$. Then the functor $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}(\mathcal{F},-):\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\sf k}\textrm{-}\mathrm{Mod})$ factors canonically through a functor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{G}))$.
Proof.
Put $X=\mathbf{G}/\mathbf{B},H=\mathbf{G}$ in [40, Appendix, Corollary A.5]
@@
Let ${\sf Inv}^{\mathbf{G}}$ denote the derived functor of invariants $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\sf k}\textrm{-}\mathrm{Mod})$.
@@
Let ${\sf F},{\sf G}\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{rep}}(\mathbf{B}))$ and $\mathcal{F}=\mathcal{L}({\sf F})$, ${\mathcal{G}}=\mathcal{L}({\sf G})\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}_{\mathbf{G}}(\mathbf{G}/\mathbf{B})$ be the associated equivariant complexes of coherent sheaves. There there is a canonical functorial isomorphism
${\sf Inv}^{\mathbf{G}}\circ\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}(\mathcal{F},\mathcal{G})\xrightarrow{\sim}\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{+}({\mathrm{Rep}}(\mathbf{B}))}({\sf F},{\sf G}).$
Proof.
Put $X=\mathbf{G}/\mathbf{B}$ and $H=\mathbf{G}$ in [40, Appendix, Proposition A.6]
The above isomorphism restricts to an isomorphism of functors ${\sf Inv}^{\mathbf{G}}\circ\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}/\mathbf{B}}(\mathcal{F},\mathcal{G})\xrightarrow{\sim}\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))}({\sf F},{\sf G}).$
8.3 $\mathbf{G}$-linear semiorthogonal decompositions
Definition 8.7.
A triangulated category $\mathcal{D}$ is called $\mathbf{G}$-linear if $\mathcal{D}$ is equipped with a monoidal action of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$, i.e. there is a bifunctor $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\times\mathcal{D}\rightarrow\mathcal{D}$. Compare [34, Section 2.7].
Our main concern is the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$ which becomes a $\mathbf{G}$-linear triangulated category under the restriction functor ${\mathrm{res}}_{\mathbf{B}}^{\mathbf{G}}:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{B}))$. Later we will also need $\mathbf{P}$, so let us use $\mathbf{P}$ instead of $\mathbf{B}$, having in mind $\mathbf{P}=\mathbf{B}$ as an important case.
Definition 8.8.
A $\mathbf{G}$-linear semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ is a collection $\langle{\sf A}_{1},\dots\sf A_{k}\rangle$ of full triangulated $\mathbf{G}$-linear subcategories such that the collection $\langle{\sf A}_{1},\dots\sf A_{k}\rangle$ is a semiorthogonal decomposition of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ in the sense of Definition 7.7.
The following proposition is just a variation of [36, Lemma 2.7]:
Proposition 8.9.
A pair of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$-linear subcategories ${\sf A},{\sf B}\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ is semiorthogonal (i.e.  ${\sf A}\subset{\sf B}^{\perp}$) if and only if the equality ${\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(N^{\ast}\otimes_{k}M)=0$ holds for all $M\in{\sf A}$, $N\in{\sf B}$.
Proof.
($\Leftarrow$). Assume given $M\in{\sf A}$, $N\in{\sf B}$ and assume ${\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(N^{\ast}\otimes_{k}M)=0$. Then
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))}(N,M)=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))}({\sf k},N^{*}\otimes M)=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}({\sf k},{\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(N^{\ast}\otimes_{k}M))=0.$
($\Rightarrow$). Let $M\in{\sf A}$, $N\in{\sf B}$ and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))}(N,M)=0$ for all such $M$, $N$. Let be $L$ an arbitrary object of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$. Then
$\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(L,{\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(N^{\ast}\otimes_{{\sf k}}M))=\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))}({\mathrm{res}}_{\mathbf{P}}^{\mathbf{G}}L,N^{\ast}\otimes_{{\sf k}}M)$
and
@@ 8.3 $\mathbf{G}$-linear semiorthogonal decompositions
the latter equality holds since the subcategory $\sf B$ is $\mathbf{G}$-linear and hence stable under tensoring with objects of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$. It follows that $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}(L,{\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(N^{\ast}\otimes_{{\sf k}}M))=0$ for an arbitrary object $L\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$; hence, ${\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(N^{\ast}\otimes_{{\sf k}}M)=0$. ∎
@@
Let us simplify notation and write $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))}$ as $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{P}}$, and $\mathop{\mathrm{RHom}}\nolimits_{\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))}$ as $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}$.
Definition 8.10.
@@
A functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(\mathbf{P}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(\mathbf{P}))$ is called $\mathbf{G}$-linear if for all $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{rep}}(\mathbf{P}))$, $N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ there are given bifunctorial isomorphisms
$\Phi({\mathrm{res}}_{\mathbf{P}}^{\mathbf{G}}(N)\otimes M)={\mathrm{res}}_{\mathbf{P}}^{\mathbf{G}}(N)\otimes\Phi(M).$
@@
    Assume given a fully faithful $\mathbf{G}$-linear functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$. Then $\Phi$ is isomorphic to the functor $\Phi_{E}(-)=(-)\otimes E$ where $E\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ such that $\mathop{\mathrm{RHom}}\nolimits_{\mathbf{P}}(E,E)={\sf k}$.
@@
    Each object $E\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$ such that ${\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(E^{\ast}\otimes E)={\sf k}$ gives a fully faithful $\mathbf{G}$-linear functor $\Phi:\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))\rightarrow\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$.
@@
    Under the assumptions of (1), the object $E$ satisfies ${\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(E^{\ast}\otimes E)={\sf k}$.
@@
(2) For $M,N\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$ we have
$\mathop{\mathrm{RHom}}\nolimits_{\mathbf{P}}(M\otimes E,N\otimes E)=\mathop{\mathrm{RHom}}\nolimits_{\mathbf{P}}(M\otimes N^{\ast},E^{\ast}\otimes E)=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}(M\otimes N^{\ast},{\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(E^{\ast}\otimes E))=\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}(M\otimes N^{\ast},{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}(M,N).$
@@
$\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}(M,{\sf k})=\mathop{\mathrm{RHom}}\nolimits_{\mathbf{P}}(M\otimes E,E)=$
$\mathop{\mathrm{RHom}}\nolimits_{\mathbf{P}}(M,E^{\ast}\otimes E)=\mathop{\mathrm{RHom}}\nolimits_{\mathbf{G}}(M,{\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(E^{\ast}\otimes E)).$
for $M\in\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{G}))$. By the Yoneda Lemma it follows that ${\mathrm{Rind}}_{\mathbf{P}}^{\mathbf{G}}(E^{\ast}\otimes E)={\sf k}$. ∎
@@
See Section 6.1 for the precise relation between the two vanishing conditions ${\rm R}^{i}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(E^{\ast}\otimes E)=0$ for $i>0$ and $\mathop{\mathrm{RHom}}\nolimits^{i}_{\mathbf{B}}(E,E)=0$ for $i>0$.
8.4 Generating $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(\mathbf{G}/\mathbf{P}))$
Definition 8.13.
@@
The set of line bundles $\langle\mathcal{L}_{\lambda}[n],\lambda\in X(\mathbf{T})_{+},n\in\mathbb{Z}\rangle$ is a generating set for $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathsf{QCoh}}(\mathbf{G}/\mathbf{P}))$.
@@
The smallest thick full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{P})$ containing $\langle\mathcal{L}_{\lambda}[n],\lambda\in X(\mathbf{T})_{+},n\in\mathbb{Z}\rangle$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{P})$.